Two identical gas containers A and B contains the masses of molecules mA = 3.34×10-27 kg and mB = 5.34×10-26 kg,
respectively. Both gases are under the same pressure and are at 10.0°C.
a) Compare the translational kinetic energy per molecule of molecules A, B and root-mean-square (rms) speeds of
them?
b) In order have both gases with the same rms speed, we need to raise the temperature of only one of these
containers. Which gas should be raised the temperature?
c) At what temperature will you accomplish your goal?
1
Expert's answer
2018-12-28T04:53:41-0500

a) At a given temperature T, all gases molecules have the same average translational kinetic energy, namely K = (3/2)kBT, where T is in Kelvin, kB = 1.38*10-23 J/K

T(K) = T(oC) + 273.15 = 10.0C + 273.15 = 283.15 K

KA = (3/2)*1.38*10-23 J/K *283.15 K = 5.86*10-21 J

KB = (3/2)*1.38*10-23 J/K *283.15 K = 5.86*10-21 J

Average translational kinetic energy of molecules A and B are the same: KA = KB = 5.86*10-21 J

υrms = sqrt(3kBT/m) (for 1 molecule)

υrms(A) = sqrt(3*1.38*10-23 J/K *283.15 K/3.34*10-27 kg) = 1873.42 m/s

υrms(B) = sqrt(3*1.38*10-23 J/K *283.15 K/5.34*10-26 kg) = 468.53 m/s

As mA < mB molecules A move faster than molecules B: υrms(A)>υrms(B).

b) υrms(A) = sqrt(3kTA/mA)

υrms (B)= sqrt(3kTB/mB)

υrms(A) = υrms (B);

sqrt(3kTA/mA) = sqrt(3kTB/mB)

3kTA/mA = 3kTB/mB

TA/mA = TB/mB

As mA < mB , then TB should be higher than TA in order υrms(A) to be equal to υrms(B). We should raise the temperature of the B container in order to have both gases with the same rms speed.

c) TA/mA = TB/mB

283.15/ 3.34*10-27 = TB / 5.34*10-26

TB = 4527 K